OpenStudy (adelyn_helps):

PLEASE HELP I WILL FAN AND MEDAL YOU What is the correct solution to this problem? A number divided by 7 is 6 less than –8. What is the number? A. –98 B. –56 C. –14 D. –7

OpenStudy (anonymous):

x/7 = -8 -6 x/7 = -14 x = -98 answer is A

OpenStudy (adelyn_helps):

k thanks

OpenStudy (anonymous):

welcome :)